# Rethinking Multimodal Learning from the Perspective of Mitigating Classification Ability Disproportion

### v2

@@ 4.1 Experimental Setup
The CREMAD dataset contains 7,442 clips, which are divided into training set with 6,698 samples and testing set with 744 samples. For KSounds dataset, which contains 19,000 video clips, is divided into training set with 15,000 clips, validation set with 1,900 clips, and testing set with 1,900 clips. VGGSound dataset includes 168,618 videos for training and validation, and 13,954 videos for testing. The NVGesture dataset is divided into 1,050 samples for training and 482 samples for testing. Twitter dataset is divided into training set with 3,197 pairs, validation set with 1,122 pairs and testing set with 1,037 pairs. Sarcasm dataset includes 19,816 pairs for the training set, 2,410 pairs for the validation set, and 2,409 pairs for the testing set. More details are provided in the appendix.
@@ 4.3 Sensitivity to Hyperparameter
Sensitivity to Threshold $\sigma$: We study the influence of threshold $\sigma$ on CREMAD dataset. The accuracy with different $\sigma\in[1,1.75]$ is shown in Figure 3. We can find that our method is not sensitive to threshold $\sigma$ in a large range.
Sensitivity to Smoothing Factor $\lambda$: We explore the influence of smoothing factor $\lambda$ on CREMAD dataset. The accuracy with different $\lambda\in[0.1,1]$ is reported in Figure 3. We can find that our method is not sensitive to hyper-parameter smoothing factor $\lambda$ in a large range.
@@ 4.4 Ablation Study
We investigate the effectiveness of our method by analyzing the influence of the key components of our objectives in Equation (2), (3), and (4), respectively denoted as $\epsilon$, $\epsilon_{o}$, and $\epsilon_{p}$. The results on CREMAD dataset are reported in Table 3. From Table 3, we can find that: (1). Both objectives in Equation (2), (3), and (4) can boost multimodal performance in terms of accuracy and MAP; (2). While the unimodal performance of the method using all objectives may not always reach the highest level, it achieves a more balanced classification performance across modalities.
@@
Table 2: The results (Acc./MAP) for ablation study on CREMAD dataset.
$\epsilon$ $\epsilon_{o}$ $\epsilon_{p}$ Multi Audio Video
0.8333/0.8967 0.6465/0.7061 0.6734/0.7523
0.8320/0.8895 0.6573/0.7235 0.6707/0.7535
0.8360/0.9014 0.6841/0.7321 0.6371/0.7335
0.8515/0.9103 0.6835/0.7529 0.6828/0.7612
@@ 5 Conclusion
This work was supported in part by the National Key R&D Program of China (2022YFF0712100), in part by the NSFC (62276131, 62506168), in part by the Natural Science Foundation of Jiangsu Province of China under Grant (BK20240081, BG2024042), and in part by the Fundamental Research Funds for the Central Universities (No.30925010205).
@@ C.2 Implementations of Toy Experiments
C.3 Influence of Unstable Confidence Score
@@
We first compare the model performance in scenario with modality noise in Table II, where $\zeta$ denotes the noise rate, “#CLS” denotes the number of classifier during training, and in the first column, we also report the fluctuation of confidence score. In fact, as training progressed, we observe that in noisy scenarios, increasing the number of classifiers beyond a certain point no longer led to improved performance on the validation set. This phenomenon became more pronounced as the noise level increased. In other words, the classifiers may overcompensate in scenarios involving modality noise. On the other hand, we also observe that after reaching its optimal performance, the model’s performance gradually degraded without a drastic drop. Therefore, in practical scenarios, an early stopping strategy based on validation set performance can be employed to select the optimal model and avoid classifier overcompensation.
Table II: The fluctuation of confidence score.
@@
Furthermore, we compare the performance with competitive baselines including naive MML, ReconBoost ReconBoost:conf/icml/CongHua24 . The results are shown in Table III. We can find that compared with the competitive method ReconBoost, our approach consistently achieves superior performance, demonstrating its effectiveness in scenario with modality noise.
Table III: Performance comparison under modality noise scenario.
@@
C.4 Computational and Memory Cost at Inference Stage
To demonstrate the practical applicability of our approach, we further provide a comprehensive analysis of its computational and memory overhead at inference stage. [^2] For computational cost during training, the analysis have been posted in Section 4.5 of the original paper. Furthermore, we conduct an experiment to analyze the computational and memory cost during inference phase on CREMAD dataset. The baselines include naive MML, PMR PMR:conf/cvpr/Fan0WW023 , AGM AGM:conf/iccv/LiLHLLZ23 , MLA MLA:conf/cvpr/ZhangYBY24 , ReconBoost ReconBoost:conf/icml/CongHua24 . The results are shown in Table IV. We can find that our method achieves superior performance while maintaining competitive inference time. Furthermore, compared with naive MML, our method introduces only 1M additional parameters when 10 classifiers are added. Given the total model size of 23.6M, this corresponds to a relatively small increase of approximately 4%.
Table IV: Computational cost comparison on CREMAD dataset.
@@ C.4 Computational and Memory Cost at Inference Stage
C.5 Selection of Score Function
@@
To some extent, entropy and loss can reflect the learning status of modality-specific models and thereby serve as proxies for their classification capabilities. To evaluate the effectiveness of using entropy and loss as score functions, we conducted experiments on the CREMAD dataset, where these two metrics were used to guide the adaptive classifier assignment. The experimental results are presented in Table V, where $\tau$ denotes the threshold. Please note that since we use the ratio between different modal metrics to compare against the threshold $\tau$, the value of $\tau$ remains consistent across different metric types such as confidence, loss, and entropy. These findings indicate that the proposed method is robust to the choice of score function, consistently achieving comparable results regardless of whether confidence score, entropy, or loss is used.
Table V: Performance with different score function.
@@
Table VI: Computational cost vs the number of classifiers.
@@
C.6 Computational Cost vs. The Number of Classifiers
The experimental results are summarized in Table VI. We observe that as the number of classifiers increases, both training time and inference time increase accordingly, while performance also improves to a certain extent. This indicates that incorporating more classifiers can indeed enhance model performance, but at the cost of increased computational overhead.

### v4

@@ 4.1 Experimental Setup
To ensure a fair comparison, we strictly follow the data partitioning strategy adopted by representative methods OGR-GB:conf/cvpr/WangTF20 ; OGM:conf/cvpr/PengWD0H22 ; MLA:conf/cvpr/ZhangYBY24 ; ReconBoost:conf/icml/CongHua24 . Specifically, the CREMAD dataset contains 7,442 clips, which are divided into training set with 6,698 samples and testing set with 744 samples. For KSounds dataset, which contains 19,000 video clips, is divided into training set with 15,000 clips, validation set with 1,900 clips, and testing set with 1,900 clips. VGGSound dataset includes 168,618 videos for training and validation, and 13,954 videos for testing. The NVGesture dataset is divided into 1,050 samples for training and 482 samples for testing. Twitter dataset is divided into training set with 3,197 pairs, validation set with 1,122 pairs and testing set with 1,037 pairs. Sarcasm dataset includes 19,816 pairs for the training set, 2,410 pairs for the validation set, and 2,409 pairs for the testing set. More details are provided in the appendix.
@@ 4.3 Sensitivity to Hyperparameter
We explore the influence of $\sigma$ and $\lambda$ on CREMAD and KSounds datasets in this section. More results can be found in appendix.
Sensitivity to Threshold $\sigma$: We study the influence of threshold $\sigma$ on CREMAD and KSounds datasets. The accuracy with different $\sigma\in[1,1.75]$ is shown in Figure 3. We can find that our method is not sensitive to threshold $\sigma$ in a large range.
Sensitivity to Smoothing Factor $\lambda$: We explore the influence of smoothing factor $\lambda$ on CREMAD and KSounds datasets. The accuracy with different $\lambda\in[0.1,1]$ is reported in Figure 3. We can find that our method is not sensitive to hyper-parameter smoothing factor $\lambda$ in a large range.
@@ 4.4 Ablation Study
We investigate the effectiveness of our method by analyzing the influence of the key components of our objectives in Equation (2), (3), and (4), respectively denoted as $\epsilon$, $\epsilon_{o}$, and $\epsilon_{p}$. The accuracy results on CREMAD and KSounds datasets are reported in Table 3. From Table 3, we can find that: (1). Both objectives in Equation (2), (3), and (4) can boost multimodal performance; (2). While the unimodal performance of the method using all objectives may not always reach the highest level, it achieves a more balanced classification performance across modalities. More results are reported in appendix.
@@
Table 2: The accuracy (Multi/Audio/Video) for ablation study on CREMAD and KSounds datasets.
$\epsilon$ $\epsilon_{o}$ $\epsilon_{p}$ CREMAD KSounds
0.8333/0.6465/0.6734 0.6942/0.5303/0.5044
0.8320/0.6573/0.6707 0.7190/0.5303/0.5145
0.8360/0.6841/0.6371 0.7198/0.5261/0.5238
0.8515/0.6835/0.6828 0.7263/0.5257/0.5736
@@ 5 Conclusion
This work was supported in part by the National Key R&D Program of China (2022YFF0712100), in part by the NSFC (62276131, 62506168), in part by the Natural Science Foundation of Jiangsu Province of China under Grant (BK20240081, BG2024042, BK20251431), and in part by the Fundamental Research Funds for the Central Universities (No.30925010205).
@@ C.2 Implementations of Toy Experiments
C.3 More Results for Ablation Study
MAP Results on CREMAD and KSounds Datasets: We have added the MAP results of the ablation study on the CREMAD and KSounds datasets in Table II. The MAP results in Table II demonstrate that the objectives defined in Equation (2), (3), and (4) all lead to performance gains.
Table II: The MAP for ablation study on CREMAD and KSounds datasets.
$\epsilon$ $\epsilon_{o}$ $\epsilon_{p}$ CREMAD KSounds
Multi Audio Video Multi Audio Video
0.8967 0.7061 0.7523 0.7796 0.5465 0.5484
0.8895 0.7235 0.7535 0.7864 0.5482 0.5606
0.9014 0.7321 0.7335 0.7821 0.5474 0.5651
0.9103 0.7529 0.7612 0.7901 0.5461 0.6040
Ablation Study on Image-Text and Tri-Modal Datasets: To further validate our method, we report an ablation study on image-text dataset, i.e., Twitter dataset, and the dataset with three modalities, i.e., NVGesture dataset. The accuracy and F1 are reported in Table III and Table IV. The results demonstrate that the objectives defined in Equation (2), (3), and (4) all lead to performance gains in terms of both accuracy and F1.
Table III: The results for ablation study on Twitter dataset.
$\epsilon$ $\epsilon_{o}$ $\epsilon_{p}$ Accuracy F1
Multi Image Text Multi Image Text
0.7425 0.5400 0.7406 0.6830 0.3957 0.6868
0.7396 0.5429 0.7387 0.6787 0.4113 0.6806
0.7445 0.5265 0.7454 0.6767 0.3865 0.6794
0.7512 0.5434 0.7488 0.6962 0.3998 0.6911
Table IV: The results for ablation study on NVGesture dataset.
$\epsilon$ $\epsilon_{o}$ $\epsilon_{p}$ Accuracy F1
Multi RGB OF Depth Multi RGB OF Depth
0.8465 0.7573 0.7780 0.7946 0.8505 0.7649 0.7809 0.7968
0.8402 0.7427 0.7718 0.7759 0.8428 0.7604 0.7720 0.7812
0.8485 0.8008 0.7988 0.7801 0.8526 0.8085 0.8009 0.7881
0.8501 0.7988 0.8029 0.7967 0.8533 0.8076 0.8057 0.8033
C.4 Evaluation of Generalization in More Scenarios
To further validate the effectiveness of our method, we employ a actor-independent data partitioning strategy on the CREMAD dataset, ensuring that the training, validation, and test sets contain mutually exclusive subjects following the setting of Emotion:conf/icassp/GoncalvesB22 ; GameTheory:conf/nips/Konstantinos25 . Specifically, we randomly select the audio-video data of 63 actors for training, 14 actors for validation, and 14 actors for testing, with no overlap of actors among the three splits.
We first investigate the influence of the hyperparameters $\sigma$ and $\lambda$ on actor-independent CREMAD dataset. The results are shown in Figure II and Figure II. The results demonstrate that our method is not sensitive to hyper-parameters $\sigma$ and $\lambda$. Furthermore, the ablation study on actor-independent CREMAD dataset is shown in Table V. The results further verify the effectiveness of the objectives defined in Equation (2), (3), and (4).
Figure I: Sensitivity to $\sigma$.
Figure II: Sensitivity to $\lambda$.
Table V: The results for ablation study on actor-independent CREMAD dataset.
$\epsilon$ $\epsilon_{o}$ $\epsilon_{p}$ Accuracy MAP
Multi Image Text Multi Image Text
0.7443 0.6197 0.5191 0.8086 0.6517 0.5433
0.7333 0.6262 0.5344 0.8135 0.6508 0.5861
0.7388 0.6273 0.5366 0.8111 0.6496 0.5860
0.7639 0.6087 0.5388 0.8300 0.6557 0.5544
Then, we compare our method with some representative methods on actor-independent CREMAD dataset. The results are shown in Table VI. From the results in Table VI, we can find that our method can achieve the best performance compared with SOTA baselines.
Table VI: Comparison with SOTA baselines on actor-independent CREMAD dataset.
Method Accuracy MAP
Audio/Video 0.6219/0.5366 0.6588/0.5380
Naive MML 0.6153 0.6671
MLA 0.7421 0.8100
ReconBoost 0.7344 0.8027
Ours 0.7639 0.8300
C.5 Influence of Unstable Confidence Score
@@
We first compare the model performance in scenario with modality noise in Table VII, where $\zeta$ denotes the noise rate, “#CLS” denotes the number of classifier during training, and in the first column, we also report the fluctuation of confidence score. In fact, as training progressed, we observe that in noisy scenarios, increasing the number of classifiers beyond a certain point no longer led to improved performance on the validation set. This phenomenon became more pronounced as the noise level increased. In other words, the classifiers may overcompensate in scenarios involving modality noise. On the other hand, we also observe that after reaching its optimal performance, the model’s performance gradually degraded without a drastic drop. Therefore, in practical scenarios, an early stopping strategy based on validation set performance can be employed to select the optimal model and avoid classifier overcompensation.
Table VII: The fluctuation of confidence score.
@@
Furthermore, we compare the performance with competitive baselines including naive MML, ReconBoost ReconBoost:conf/icml/CongHua24 . The results are shown in Table VIII. We can find that compared with the competitive method ReconBoost, our approach consistently achieves superior performance, demonstrating its effectiveness in scenario with modality noise.
Table VIII: Performance comparison under modality noise scenario.
@@
C.6 Computational and Memory Cost at Inference Stage
To demonstrate the practical applicability of our approach, we further provide a comprehensive analysis of its computational and memory overhead at inference stage. [^2] For computational cost during training, the analysis have been posted in Section 4.5 of the original paper. Furthermore, we conduct an experiment to analyze the computational and memory cost during inference phase on CREMAD dataset. The baselines include naive MML, PMR PMR:conf/cvpr/Fan0WW023 , AGM AGM:conf/iccv/LiLHLLZ23 , MLA MLA:conf/cvpr/ZhangYBY24 , ReconBoost ReconBoost:conf/icml/CongHua24 . The results are shown in Table IX. We can find that our method achieves superior performance while maintaining competitive inference time. Furthermore, compared with naive MML, our method introduces only 1M additional parameters when 10 classifiers are added. Given the total model size of 23.6M, this corresponds to a relatively small increase of approximately 4%.
Table IX: Computational cost comparison on CREMAD dataset.
@@ C.6 Computational and Memory Cost at Inference Stage
C.7 Selection of Score Function
@@
To some extent, entropy and loss can reflect the learning status of modality-specific models and thereby serve as proxies for their classification capabilities. To evaluate the effectiveness of using entropy and loss as score functions, we conducted experiments on the CREMAD dataset, where these two metrics were used to guide the adaptive classifier assignment. The experimental results are presented in Table X, where $\tau$ denotes the threshold. Please note that since we use the ratio between different modal metrics to compare against the threshold $\tau$, the value of $\tau$ remains consistent across different metric types such as confidence, loss, and entropy. These findings indicate that the proposed method is robust to the choice of score function, consistently achieving comparable results regardless of whether confidence score, entropy, or loss is used.
Table X: Performance with different score function.
@@
C.8 Computational Cost vs. The Number of Classifiers
The experimental results are summarized in Table XI. We observe that as the number of classifiers increases, both training time and inference time increase accordingly, while performance also improves to a certain extent. This indicates that incorporating more classifiers can indeed enhance model performance, but at the cost of increased computational overhead.
Table XI: Computational cost vs the number of classifiers.